OpenStudy (anonymous):

Expand (5+z)square root 4

jagr2713 (jagr2713):



jagr2713 (jagr2713):

like that?

OpenStudy (anonymous):

yes

jagr2713 (jagr2713):

5+2=7 7*7*7*7=2401

OpenStudy (anonymous):

no the letter z

jagr2713 (jagr2713):

oihohoh i thought it was a 2

jagr2713 (jagr2713):

625

jagr2713 (jagr2713):

5*5*5*5=625

OpenStudy (anonymous):

keep going 

jagr2713 (jagr2713):

thats it

OpenStudy (anonymous):

625+250z+100z+10z+z 625+500z+150z+20z+z 625+375z+150z+5z+z 625+125z+25z+5z+z like this but I need the correct one